OpenStudy (anonymous):

Solve the equation by completing the square: 3x2 - 24x = 27. Enter the largest value of the coordinates.

OpenStudy (anonymous):

assuming this is an on line course (enter the largest value is a hint) then just find the zeros without completing the square if you want to know how to complete the square, that is a different problem

OpenStudy (nenadmatematika):

3x^2-24x-27=0 x^2-8x-9=0 (x-9)(x+1)=0 x=9 or x=-1 so the answer is 9

OpenStudy (anonymous):

start with \[x^2-8x-9=0\] \[(x+1)(x-9)=0\] \[x=-1,x=9\]